Ratio of 5 paise to 25 paise is the same as the ratio of 20 paise to _____

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Ratio of 5 paise to 25 paise is the same as the ratio of 20 paise to 100 paise. From the question,

5 paise : 25 paise = 20 paise: [ ]
Let us assume the missing number be Q, 5 paise : 25 paise = 20 paise: Q
5/25 = 20/Q
Q = (20 × 25)/5
Q = 500/5
Q = 100
Therefore, Ratio of 5 paise to 25 paise is the same as the ratio of 20 paise to 100 paise

Thanks